OpenStudy (anonymous):

Concrete can be purchased by the cubic yard. How much will it cost to pour a slab 18 feet by 18 feet by 6 inches for a patio if the concrete costs $42.00 per cubic yard? $252.00 $756.00 $1,134.00 $6,804.00

OpenStudy (welshfella):

How many cubic feet are there in a cubic yard

OpenStudy (durply):

Hi

OpenStudy (welshfella):

there are 3*3*3 sq ft in a sq yd

OpenStudy (welshfella):

volume of the slab = 18*18*0.5 cu ft

OpenStudy (welshfella):

I mean 3*3*3 = 27 cu ft not sq ft

OpenStudy (welshfella):

now convert the volume of the slab to cu yds by dividing it by 27

OpenStudy (welshfella):

the cost will be 42 * this volume

OpenStudy (welshfella):

volume of the slab in cu yds = 18*18*0.5 -------- 27

OpenStudy (welshfella):

grab your calculator and work that out then multiply the result by 42 to find the required cost

OpenStudy (anonymous):

ok got it thnx!

OpenStudy (welshfella):

yw